In a small country, the demand curve is given as: Q=100-5P, supply curve: Q=3P-12, and the world price is $10.

I. What is the social surplus under free trade? (5%)

II. If the government impose a $2/unit tariff on the good, what is the deadweight loss? (10%)

III. Show the change in equilibrium and deadweight loss on a graph. (10%)
Answer #1

I. Consumer surplus=(1/2*(20-10)*50) = $250

Producer surplus= 1/2*(10-4)*18= $54

Social surplus= $250+$54= $304

II. Deadweight loss= 1/2*(2)(24-18)+(1/2*2*(50-40))= 6+10= $16

III. Equilibrium changes from closed economy equilibrium E1 to free market equilibrium E2 to after tariff equilibrium E3. Price Domestic Supply -16 -14 12 DWL DWL 10 Demand 00 5 10 15 20 25 30 35 40 45 50 55 Quantity